OpenStudy (falling_in_katt):

WILL MEDAL Match A.If b = 2 and c = 3, then a = B.If a = 4 and b = 6, then c = C.If a = 3 and c = 6, then b =

OpenStudy (anonymous):

is that the whole question?

OpenStudy (falling_in_katt):

match with 1.\[2\sqrt{13}\] 2.\[\sqrt{5}\] 3.\[3\sqrt{3}\]

OpenStudy (anonymous):

i think its A3 B1 C2

OpenStudy (anonymous):

tell me how it goes

OpenStudy (falling_in_katt):

It went good thank you

OpenStudy (anonymous):

np